OpenStudy (anonymous):

What is the solution of the equation 4x + 17 = 45?

OpenStudy (january123):

isolate x by first subtracting 17 from both side

OpenStudy (anonymous):

4x + 17= 45 - 17 -17 4x = 28 /4 /4 x = 7

OpenStudy (anonymous):

Hi! I hope I can help you today. :) First, subtract 17 from both sides. {4x=45-17} Second, simplify 45-17 (to get 28)} {4x=28} Divide both sides by 4 {x=28/4} And simplify 28/4 to 7. *x = 7*

OpenStudy (anonymous):

thanks for the help you guys